# Generative AI for Cel-Animation: A Survey

### v3

@@ 2.2.2 GAN, VAE, and Diffusion Models
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=6em, edge+=darkgray, line width=1pt, s sep=20pt, inner xsep=3pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.0em,font=,, where level=2text width=10em,font=,, where level=3text width=8.2em,font=,, where level=4text width=8.2em,font=,, [ GenAI for
@@
3)[ Pre-production
3.1) [ Scripting (§3.1) [ HoLLMwood [14], ChatGPT [1], Gemini [86], Claude [4],
LLaMA [4], InternVL [18], Qwen [5] , leaf, text width=26em ] ] [ Setting (§3.1) [ Midjourney [79], Stable Diffusion [87] , leaf, text width=17em ] ] [ Storyboarding (§3.1) [ StoryWeaver [143], CogCartoon [157], Coherent Story [84],
Seed-Story [138], Make-A-Story [85], Make-A-Storyboard [107],
Storyboarder.ai [104], Animate-A-Story [37], StoryImager [112],
VideoDirectorGPT [64], CustomCrafter [130], VideoStudio [73],
Mind the Time [131], DREAMRUNNER [123], Vlogger [158],
Anim-Director [62], StoryGen [66], , leaf, text width=28em ] ] ] [ Production
3.2) [ Layout (L/O) (§3.2) [ CogCartoon [157], SGSIG [148], VideoComposer [122],
LayoutGAN [60], Layout2Vid [64], DiffSensei [129],
MangaDiffusion [16], CameraCtrl [36] , leaf, text width=24em ] ] [ Keyframe
Animation
 (§3.2) [ AnimateAnyone [39], Champ [156], MimicMotion [149],
Animate-X [108], MikuDance [144], CoNR [65] , leaf, text width=24.5em ] ] [ Inbetweening (§3.2) [ ToonCraft [132], JointStroke [80], AniSora [49], AnimeInterp [96],
AnimeInbet [30], AutoFI [92], SAIN [91], MotionBridge [111],
toona.io [116], DeepSketch-Guided [61], DeepGeometrized [97],
LayerAnimate [140] , leaf, text width=28.5em ] ] [ Colorization (§3.2) [ AniDoc [78], ToonCraft [132], VToonify [135], TokenFlow [32],
StyleGANEX [136], FRESCO [139], toona.io [116], TRE-Net [120],
LayerAnimate [140], DeepLineArt [95], SketchBetween [72],
Flow Estimate-Refine Net [141], PromptFix [142], MangaNinja [71] ,
Animation Xformer [10], Inclusion Matching [28], LVCD [45] , leaf, text width=29em ] ] ] [ Post-production
3.3) [ Compositing &
Photography
 (§3.3) [ IC-Light [3], Dovenet [24], High res harmonization [25],
Pct-net [35], SSH [48], Object Compositing [113], SSN [93],
Dr.Bokeh [94], Floating No More [75], ObjectDrop [127],
Alchemist [88], Disenstudio [13], LayerAnimate [140] , leaf, text width=27.5em ] ] [ Cutting (CT) (§3.3) [ M-SAN [109], Reframe Anything [8], OpusClip [83] , leaf, text width=24em ] ] [ Music & Sound
Effects
 (§3.3) [ FoleyMusic [31], Video2Music [51], V2Meow [106],
MelFusion [20], VidMuse [114], SpecVQGAN [46],
Im2Wav [89], FoleyGen [76], DiffFoley [74],
Action2Sound [12] , leaf, text width=24em ] ] [ After Record (AR) &
Dubbing (DB)
 (§3.3) [ StyleDubber [23], ANIM-400K [7], EmoDubber [22],
MovieDubbing [150], HPMDubbing [21], V2C [15] , leaf, text width=25em ] ] ] [ Others (§3.4) [ Cel-Animation
Editing (§3.4) [ Re:Draw [9], ScalingConcept [42] , leaf, text width=17em ] ] [ Cels Decomposition
3.4) [ Sprite-from-Sprite [145], Generative Omnimatte [57],
TransPixar [119], LayerAnimate [140] , leaf, text width=23em ] ] [ 3D Assistance (§3.4) [ Toon3D [124], CoNR [65], DrawingSpinUp [153] , leaf, text width=23em ] ] ] ]

### v4

@@ 2.2.2 GAN, VAE, and Diffusion Models
[…]
GenAI for
@@
3)
Pre-production 3.1) Scripting (§3.1) HoLLMwood [14], ChatGPT [1], Gemini [86], Claude [4], LLaMA [4], InternVL [18], Qwen [5] Setting (§3.1) Midjourney [79], Stable Diffusion [87] Storyboarding (§3.1) StoryWeaver [143], CogCartoon [157], Coherent Story [84], Seed-Story [138], Make-A-Story [85], Make-A-Storyboard [107], Storyboarder.ai [104], Animate-A-Story [37], StoryImager [112], VideoDirectorGPT [64], CustomCrafter [130], VideoStudio [73], Mind the Time [131], DREAMRUNNER [123], Vlogger [158], Anim-Director [62], StoryGen [66], Production 3.2) Layout (L/O) (§3.2) CogCartoon [157], SGSIG [148], VideoComposer [122], LayoutGAN [60], Layout2Vid [64], DiffSensei [129], MangaDiffusion [16], CameraCtrl [36] Keyframe Animation (§3.2) AnimateAnyone [39], Champ [156], MimicMotion [149], Animate-X [108], MikuDance [144], CoNR [65] Inbetweening (§3.2) ToonCraft [132], JointStroke [80], AniSora [49], AnimeInterp [96], AnimeInbet [30], AutoFI [92], SAIN [91], MotionBridge [111], toona.io [116], DeepSketch-Guided [61], DeepGeometrized [97], LayerAnimate [140] Colorization (§3.2) AniDoc [78], ToonCraft [132], VToonify [135], TokenFlow [32], StyleGANEX [136], FRESCO [139], toona.io [116], TRE-Net [120], LayerAnimate [140], DeepLineArt [95], SketchBetween [72], Flow Estimate-Refine Net [141], PromptFix [142], MangaNinja [71] , Animation Xformer [10], Inclusion Matching [28], LVCD [45] Post-production 3.3) Compositing & Photography (§3.3) IC-Light [3], Dovenet [24], High res harmonization [25], Pct-net [35], SSH [48], Object Compositing [113], SSN [93], Dr.Bokeh [94], Floating No More [75], ObjectDrop [127], Alchemist [88], Disenstudio [13], LayerAnimate [140] Cutting (CT) (§3.3) M-SAN [109], Reframe Anything [8], OpusClip [83] Music & Sound Effects (§3.3) FoleyMusic [31], Video2Music [51], V2Meow [106], MelFusion [20], VidMuse [114], SpecVQGAN [46], Im2Wav [89], FoleyGen [76], DiffFoley [74], Action2Sound [12] After Record (AR) & Dubbing (DB) (§3.3) StyleDubber [23], ANIM-400K [7], EmoDubber [22], MovieDubbing [150], HPMDubbing [21], V2C [15] Others (§3.4) Cel-Animation Editing (§3.4) Re:Draw [9], ScalingConcept [42] Cels Decomposition 3.4) Sprite-from-Sprite [145], Generative Omnimatte [57], TransPixar [119], LayerAnimate [140] 3D Assistance (§3.4) Toon3D [124], CoNR [65], DrawingSpinUp [153]